# Oral exams in introductory statistics class with non-native English speakers

### v3

@@ 1 Introduction
Oral exams are a unique tool to assess student’s understanding of course material. This is especially prescient in introductory statistics courses where many topics are conceptually difficult, e.g., interpretation of probability, $p$-values, hypothesis testing, etc. These challenges are only heightened for students learning in a second language. Specifically, in this work we are interested in an introductory statistics course of Japanese students learning in English. In this context, oral exams provide many benefits for assessing understanding. First, they give instructors a clearer picture of student’s conceptual mastery (Theobold,, 2021; Sabin et al.,, 2021). As discussed in the Guidelines for Assessment and Instruction in Statistics Education (GAISE), conceptual understanding is a fundamental goal for statistics educators (Carver et al.,, 2016). Specifically, this means instructors should “view the primary goal as to discover and apply concepts” and “focus on students’ understanding of key concepts, illustrated by a few techniques, rather than covering a multitude of techniques with minimal focus on underlying ideas” (Carver et al.,, 2016). Unfortunately, far too many introductory statistics courses focus on formulas and memorization, with little emphasis on conceptual understanding. This is even more so the case in the Japanese context, where memorization is a major pillar of primary through higher education (e.g., Nemoto,, 1999). In particular, Dunn, (2015) highlights a lack of critical thinking skills in Japanese college freshmen, largely stemming from a secondary education pedagogical framework which stresses memorization. This poses a major challenge if the goal of an introductory statistics course should be to ensure students are “able to make decisions about the most appropriate ways to visualize, explore, and, ultimately, analyze a set of data” Carver et al., (2016).
@@
The importance of assessing mastery has only increased with the current proliferation of large language models (LLMs), e.g., ChatGPT, Gemini, etc. It is now easier than ever for students to circumvent the learning process by simply passing off their assignments to LLMs (Paustian and Slinger,, 2024; Al Labadi and Ly,, 2025). This problem is pervasive for students learning in a second language where the stress of deadlines and writing in another language may push them to use an LLM, or assignments may be completed in Japanese only to have an application translate it to English. It can be hard to catch such cases, as LLM checkers have been shown to be biased against non-native English-speakers (Liang et al.,, 2023). Specifically, the authors show that such checkers are more likely to incorrectly flag responses written by non-native speakers as coming from an LLM when in fact it was genuinely produced by a student, when compared with native speakers. Ideally, with the prospect of oral exams, students should be motivated to truly learn the material (Delson et al.,, 2022) instead of inappropriately relying on LLMs. Moreover, not only do oral exams necessitate deep, conceptual understanding of statistical concepts, but they also allow students to build their oral communication skills (Hazen,, 2020; Sabin et al.,, 2021).
@@ 2.1 University
Oral exams were given to an introductory statistics class of nine students at Akita International University (AIU) in Akita, Japan during the spring 2025 semester. AIU is well-known in Japan for its unique learning environment; while over 90% of degree-seeking students are Japanese, classes are taught entirely in English. Students are admitted with fairly strong English skills and then spend their first year honing these skills to an academic level. Furthermore, AIU is modeled after an American liberal arts college. While many Japanese universities have become highly specialized or technical, AIU seeks to form well-rounded, globally-minded individuals.
@@ 2.2 Course details
Lastly, ideas were illustrated using R in almost every lesson. The course topics generally followed Learning Statistics with R (Navarro,, 2013), a textbook that integrates coding with concepts and is aimed at non-statistics majors. Beyond hoping that students would gain some programming skills, the main motivation for incorporating R was to aid in the understanding of certain concepts. Sampling distributions, the Law of Large Numbers and the Central Limit Theorem are just some of the concepts that introductory students typically struggle with. By using R to “see” these concepts play out, I hoped that students would gain a deeper appreciation. That being said, the midterm exam did not consist of any questions about R.
@@ 2.3 Previous implementations
This was my third time conducting oral exams in this course as I had administered them during the previous two semesters. These experiences paved the way for the current implementation. In the first iteration in spring 2024, students were given an oral midterm and final exam. The midterm was ten minutes and contained two definition questions and two conceptual questions. The final exam was thirty minutes and included three questions analyzing the results of a model or test. I found the final to be extremely long, and even so, many students could not answer all of the questions in the allotted time. Thus, in the fall 2024 semester, the final exam was converted to a written exam, but the midterm format was kept roughly the same. During this second implementation, I realized that four questions was too much for only ten minutes. In particular, there was not enough time to probe student’s answer as I needed to rush to the next question. Additionally, test performance was generally quite poor. Based on these experiences, the test was modified to its current state, described below.
@@ 3.1 Format
During the exam, I graded and took notes while also recording the exams. The purpose of the recording was to ensure fair grading across students, e.g., the same answers by different students received the same grade, and as an insurance policy in-case students disputed their grade and/or answers. Indeed, in a previous implementation when I did not record the exams, a student protested part of their grade but since there was no recording, it was impossible to adjust. After the exam, I also gave some brief feedback to each student on their performance. Finally, the study received Institutional Review Board (IRB) approval from my university.
@@ 3.2 Questions and scoring
Approximately one week before the exam, students were told some topics that might appear on the exam. All exam questions came from this list and each student was given the same exam. Each question was scored out of 4 points (with a granularity of 0.5 points). After scoring, I compared the results with the mastery rubric from Theobold, (2021) and the scores generally tended to agree. The exam questions are below with italicized writing briefly explaining the grading.
@@ 3.3 Results
The exam performance for each student is reported in Figure 1. The average score (out of twelve) was 9.9 with a standard deviation of 1.1. The median and mode were both 9.5. Separated by individual question, the average (out of four) and standard deviation for the $p$-value, probability interpretation and Central Limit Theorem questions were 2.9 (0.7), 3.8 (0.3) and 3.2 (0.6), respectively.
In Figure 2, the student’s oral exam score and TOEFL iTP score are plotted for the seven students who reported this value, along with the least squares regression line. The correlation coefficient is $-0.65$. Based on their self-assessed speaking ability, four students said they had “weak” or “very weak” speaking skills, while the remaining five choose “intermediate” or “fluent.” The students in the former group had a mean score of 9.5, while the mean of the latter group was 10.2. Finally, three of the top four scores were from students with typical English ability compared to others at the university, based on my subjective assessment of the student’s language ability.
Figure 1: Dot plot of exam performance where each dot represents a student’s score.
Figure 2: Scatter plot of exam performance against TOEFL iTP scores.
@@ 3.5 Student Feedback
Several students said, however, that they thought the oral exam helped them to identify weak points in their understanding. For example: “after finishing the test, I thought [the] oral test is good to make sure how I understand what I learn because it was good chance to output my knowledge” and, “the oral exam clarifies what I understand and what I do not understand.” Students also liked that they could easily ask if they did not understand the question: “one thing good [sic] about [the] oral exam is that I can ask [the] professor when I cannot fully understand the meaning or purpose of the question. Also, professor ask[ed] me further explanation to confirm my understanding, and then I can describe more to tell my interpretation.”
@@
3.6 Discussion
First, the scores were generally much higher than in previous semesters. This is likely because I gave more specific information about what to expect on the exam, i.e., example questions. Additionally, I may have graded more leniently since I was calibrated as to the expected responses. Next, the most difficult questions were: “why we need to assume the null hypothesis is true to calculate the $p$-value?” and, “what is the mean and variance of the normal distribution in the Central Limit Theorem?” These were the most mathematical questions on the exam, where student’s tend to have the greatest weakness. Moreover, since I told them there would not be calculations on the exam, they have have supposed that these types of questions were also off-limits.
One of the main advantages of oral exams is the ability to probe unclear answer to truly assess student’s understanding. An example of this was detailed in the previous sub-section with the Central Limit Theorem question. In general, when I probed I was looking for students to use the precise language we discussed in-class, but as the probability response above showed, sometimes a student’s organic response demonstrated a deeper understanding than my expectation. An unforeseen benefit of the exams was that this also works in the opposite direction; students can ask for clarification when they do not understand the question, as noted in the student feedback. Indeed, I notice that in many written questions both in this course and others, student’s often misunderstand the question, so this was an unexpected bonus of using oral exams, and aligned with Ohmann, (2019).
As for the main interest in this paper, the analysis above shows almost no relationship between English proficiency and test scores. For those who reported their TOEFL iTP score, the higher scoring students on this standardized test actually performed slightly worse on the oral exam. Also, only one of the top four scores came from a student with above-average English ability compared to his/her peers. Additionally, based on student’s self-perception of their English abilities, there was less than a one point difference in the average score for those who were more confident in their ability. One surprising finding was that the better English speakers tended to get more flustered when they did not completely understand a question. For example, in the probability interpretation question, a few students were particularly confused by the phrase “what would you say to your friend?” It is unclear if this negatively affected these students, but their reaction was notably different from the other students. I hypothesize that the weaker English students may not have fully understood the question and resorted to answering about the topic more generally, while the stronger English students wanted to precisely answer the question. Even though students felt that they were hindered from sharing their knowledge in English, the general trends of these results encouraged me that I was truly testing their statistical understanding, and not merely their English abilities.
@@ 4 Conclusion
In this work, I describe my experience giving oral exams in an introductory statistics class of non-native English speakers at a Japanese university. The particular focus is on the effects of students learning in their second language and, more specifically, if students with stronger English skills had an advantage on the exam. Based on my experience administering the exam, as well as the grade outcomes, there was not a strong relationship between students’ performance and English ability. While the sample size was fairly small, the results agreed with what I had seen in previous semesters as well.
@@
Overall, I was fairly pleased with the results of the oral exams. I would give several points of consideration to instructors wishing to implement similar exams, especially in a context with non-native English speakers. First, I encourage giving similar practice problems as well as sufficient hints about the exam contents. Even with repeated statements about the exam material, many students still could not clearly articulate a deep understanding of these concepts, e.g., why we assume the null hypothesis is true. In the non-native English context where listening and speaking comprehension skills are lower, it is vital to give sufficient reminders.
@@
When working with non-native speakers, I also found it imperative to give an adequate amount of time for the exam. I thought that fifteen minutes would be enough based on my previous implementations, but it was still insufficient for many students. As non-native speakers tend to take more time to understand the question, including requiring it to be read slowly and multiple times, it is important to plan sufficient time. Indeed, several student were pressed for time in a way that meant I could not ask all the probing follow-up questions that I would have liked, especially on the last question.
@@
I also found the grading to be more difficult than I expected, particularly for partially correct answers. For example, students might have said an extreme result was more “surprising” but could not clearly articulate what this meant. Moreover, it was challenging to keep consistent grading expectations throughout the day, when the first and last exam were close to eight hours apart. For example, I started to realize that one of the Central Limit Theorem follow up questions was particularly confusing for students, so half-way though the day, I changed what I was looking for in this response. Since I could re-listen to the previous responses and adjust their grades accordingly, this was not a problem, but still it created some challenges for fairness. If you plan to give an oral exam, the more precise and granular the rubric you can make, the better.
@@
In summary, I was encouraged by the results of the exams. The test did not merely gauge a student’s English ability, but yielded a meaningful picture of their statistical understanding. Even while working in their second-language, I was impressed by many of the students ability to express complex statistical concepts. Moreover, the exam setting yielded the added benefit of being able to immediately encourage students who did particularly well, easing much of their stress as they left my office.
@@ Acknowledgments
I would like to thank Sunghwan Byun and Herle McGowan for their helpful comments. I also thank the two anonymous referees and editor whose comments greatly improved the quality of the manuscript.

### v5

@@ 1 Introduction
Oral exams are a unique tool to assess student’s understanding of course material. This is especially relevant in introductory statistics courses where many topics are conceptually difficult, e.g., interpretation of probability, $p$-values, hypothesis testing, etc. These challenges are only heightened for students learning in a second language. Specifically, in this work we are interested in an introductory statistics course of predominantly Japanese students learning in English. In this context, oral exams provide many benefits for assessing understanding. First, they give instructors a clearer picture of students’ conceptual mastery (Theobold,, 2021; Sabin et al.,, 2021). As discussed in the Guidelines for Assessment and Instruction in Statistics Education (GAISE), conceptual understanding is a fundamental goal for statistics educators (Carver et al.,, 2016). Specifically, this means instructors should “view the primary goal as to discover and apply concepts” and “focus on students’ understanding of key concepts, illustrated by a few techniques, rather than covering a multitude of techniques with minimal focus on underlying ideas” (Carver et al.,, 2016). Unfortunately, far too many introductory statistics courses focus on formulas and memorization, with little emphasis on conceptual understanding. This is even more so the case in the Japanese context, where memorization is a major pillar of primary through higher education (e.g., Nemoto,, 1999). In particular, Dunn, (2015) highlights a lack of critical thinking skills in Japanese college freshmen, largely stemming from a secondary education pedagogical framework which stresses memorization. This poses a major challenge if the goal of an introductory statistics course should be to ensure students are “able to make decisions about the most appropriate ways to visualize, explore, and, ultimately, analyze a set of data” (Carver et al.,, 2016).
@@
The importance of assessing mastery has only increased with the current proliferation of large language models (LLMs), e.g., ChatGPT, Gemini, etc. It is now easier than ever for students to circumvent the learning process by simply passing off their assignments to LLMs (Paustian and Slinger,, 2024; Al Labadi and Ly,, 2025). This problem is pervasive for students learning in a second language where the stress of deadlines and writing in another language may push them to use an LLM, or assignments may be completed in Japanese only to have an application translate it to English. Ideally, with the prospect of oral exams, students should be motivated to truly learn the material (Delson et al.,, 2022) instead of inappropriately relying on LLMs. Moreover, not only do oral exams necessitate deep, conceptual understanding of statistical concepts, but they also allow students to build their oral communication skills (Hazen,, 2020; Sabin et al.,, 2021).
@@ 2.1 University
Oral exams were given to an introductory statistics class of nine students at Akita International University (AIU) in Akita, Japan during the spring 2025 semester. AIU is well-known in Japan for its unique learning environment; while the vast majority of degree-seeking students are Japanese, classes are taught entirely in English. Students are admitted with fairly strong English skills and then spend their first year honing these skills to an academic level. Specifically, students must complete the English for Academic Purposes (EAP) program to improve their reading, writing, speaking and listening skills. These courses are designed to help students improve their English proficiency such that they are prepared to learn (other subject matters) in English. Furthermore, AIU is modeled after an American liberal arts college. While many Japanese universities have become highly specialized or technical, AIU seeks to form well-rounded, globally-minded individuals.
@@ 2.2 Course details
Lastly, ideas were illustrated using R for almost every topic. The course topics generally followed Learning Statistics with R (Navarro,, 2013), a textbook that integrates coding with concepts and is aimed at non-statistics majors. Beyond hoping that students would gain some programming skills, the main motivation for incorporating R was to aid in the understanding of certain concepts. Sampling distributions, the Law of Large Numbers and the Central Limit Theorem are just some of the concepts that introductory students typically struggle with. By using R to “see” these concepts play out, I hoped that students would gain a deeper appreciation. That being said, the midterm exam did not consist of any questions about R.
@@ 2.3 Previous implementations
This was my third time conducting oral exams in this course as I had administered them during the previous two semesters. These experiences paved the way for the current implementation. In the first iteration in spring 2024, students were given an oral midterm and final exam. The midterm was ten minutes and contained two definition questions and two conceptual questions. The final exam was thirty minutes and included three questions analyzing the results of a model or test. I found the final to be extremely long, and even so, many students could not answer all of the questions in the allotted time. Thus, in the fall 2024 semester, the final exam was converted to a written exam, but the midterm format was kept roughly the same. During this second implementation, I realized that four questions was too much for only ten minutes. In particular, there was not enough time to probe student’s answer as I needed to rush to the next question. Additionally, test performance was generally quite poor. Finally, there were multiple versions of the exam in these previous implementations, and I did not record them. Based on these experiences, the test was modified to its current iteration, described below.
@@ 3.1 Format
During the exam, I graded and took notes while also recording the exam with a smartphone application. The purpose of the recording was to ensure fair grading across students, e.g., the same answers by different students received the same grade, and as an insurance policy in case students disputed their grade and/or answers. Indeed, in a previous implementation when I did not record the exams, a student protested part of their grade but since there was no recording, it was impossible to adjust. After the exam, I also gave some brief feedback to each student on their performance. This study received Institutional Review Board (IRB) approval from my university. Lastly, the final exam in the course was a more traditional written exam.
@@ 3.2 Questions and scoring
Approximately one week before the exam, students were told some topics that might appear on the exam. All exam questions came from this list and each student was given the same exam. Each question was scored out of 4 points (with a granularity of 0.5 points). After scoring, I compared the results with the mastery rubric from Theobold, (2021) (included in the Appendix) and the scores generally tended to agree. As the only instructor for the course, achieving consistent and reliable grading was difficult, but recording and re-listening to the exams helped mitigate these challenges. The exam questions are below with italicized writing briefly explaining the grading.
@@ 3.3 Results
The oral midterm exam performance for each student is reported in Figure 1 along with their self-assessed speaking ability. The average score (out of twelve points) was $9.9$ with a standard deviation of $1.1$. The median and mode were both $9.5$. Separated by individual question, the average (out of four) and standard deviation for the $p$-value, probability interpretation and Central Limit Theorem questions were $2.9$ ($0.7$), $3.8$ ($0.3$) and $3.2$ ($0.6$), respectively. Additionally, the correlation between the oral midterm and (written) final exam scores was $0.80$.
In Figure 2, the student’s oral exam score and TOEFL iTP score (out of $203$ points) are plotted for the seven students who reported this value. Based on their self-assessed speaking ability, four students said they had “weak” or “very weak” speaking skills, while the remaining five choose “intermediate” or “fluent.” The students in the former group had a mean score of $9.5$, while the mean of the latter group was $10.2$. On the final exam, these means were $78$ and $82$ (out of $100$), respectively. Finally, three of the top four scores were from students with typical English ability compared to others at the university, based on my subjective assessment of the student’s language ability.
Figure 1: Scatter plot of oral midterm exam performance (out of twelve points) against self-assessed speaking ability.
Figure 2: Scatter plot of oral midterm exam performance (out of twelve points) against TOEFL iTP score (out of $203$ points).
@@ 3.5 Student Feedback
Several students said, however, that they thought the oral exam helped them to identify weak points in their understanding. For example: “after finishing the test, I thought [the] oral test is good to make sure how I understand what I learn because it was good chance to output my knowledge” and, “the oral exam clarifies what I understand and what I do not understand.” Students also liked that they could easily ask if they did not understand the question: “one thing good [sic] about [the] oral exam is that I can ask [the] professor when I cannot fully understand the meaning or purpose of the question. Also, [the] professor ask[ed] me further explanation to confirm my understanding, and then I can describe more to tell my interpretation.”
@@ 4.1 Main findings
In this work, I describe my experience giving oral exams in an introductory statistics class of non-native English speakers at a Japanese university. The particular focus is on the effects of students learning in their second language and, more specifically, if students with stronger English skills had an advantage on the exam. There were several major takeaways from this experience.
First, the scores were generally much higher than in previous semesters. This is likely because I gave more specific information about what to expect on the exam, i.e., example questions. Additionally, I may have graded more leniently since I was calibrated to the expected responses. Next, the most difficult questions were: “why we need to assume the null hypothesis is true to calculate the $p$-value?” and, “what is the mean and variance of the normal distribution in the Central Limit Theorem?” Since I told students that there would not be calculations on the exam, they may have have supposed that these types of questions were also off-limits, leading to the lower performance.
One of the main advantages of oral exams is the ability to probe unclear answers to truly assess student’s understanding. An example of this was detailed in the previous sub-section with the Central Limit Theorem question. In general, when I probed, I was looking for students to use the precise language we discussed in-class, but as the probability response showed (contrasting the individual and government’s perspective), sometimes a student’s organic response demonstrated a deeper understanding than my expectation. An unforeseen benefit of the exams was that the ability to probe unclear responses also works in the opposite direction; students can ask for clarification when they do not understand the question, as noted in the student feedback. Indeed, I noticed that in many written questions, both in this course and others, students often misunderstand the question, so this was an unexpected bonus of using oral exams, and aligned with Ohmann, (2019) who also found students benefiting from the conversation style of the exam.
As for the main research question, the results of Section 3.3 show almost no relationship between English proficiency and test scores. For those who reported their TOEFL iTP score, the higher scoring students on this standardized test actually performed slightly worse on the oral exam. Also, one of the top four scores came from a student with above-average English ability compared to his/her peers. Additionally, based on student’s self-perception of their English abilities, those who were more confident in their ability scored $0.7$ points ($6\%$) higher on average. A comparable difference of $4\%$ was found on the final (written) exam as well. Moreover, the high correlation between the written and oral exam grades further supports the claim that the oral exam was testing statistical ability and not English speaking proficiency.
One surprising finding was that the better English speakers tended to get more flustered when they did not completely understand a question. For example, in the probability interpretation question, a few students were particularly confused by the phrase “what would you say to your friend?” It is unclear if this negatively affected these students, but their reaction was notably different from the other students. I hypothesize that the weaker English students may not have fully understood the question and resorted to answering about the topic more generally, while the stronger English students wanted to precisely answer the question. In summary, these results seem to indicate that oral exams are a fair assessment tool even with students of varied English abilities. In other words, oral exams appear to truly test statistical understanding instead of merely English proficiency, and this should give instructors the confidence to experiment with them in their own courses.
These findings are an important contribution to the English-medium instruction (EMI) literature. For example, many works assert that EMI can be detrimental to student’s content learning, but there has been minimal quantitative proof (Macaro et al.,, 2018). Here, we showed that there does not appear to be a large effect on student’s content learning based on their English proficiency. We acknowledge, however, that the instructor was a native English speaker, so there may be added difficulties if that is not the case (e.g., Tange,, 2010; Bradford,, 2016).
4.2 Recommendations and limitations
@@
Overall, I was fairly pleased with the results of the oral exams. I would give several points of consideration to instructors wishing to implement similar exams, especially in a context with non-native English speakers. First, I encourage giving similar practice problems and low-stakes practice opportunities, as well as sufficient hints about the exam content. Even with repeated statements about the exam material, many students still could not clearly articulate a deep understanding of these concepts, e.g., why we assume the null hypothesis is true. In the non-native English context where listening and speaking comprehension skills are lower, it is vital to give sufficient reminders.
@@
When working with non-native speakers, I also found it imperative to give an adequate amount of time for the exam. I thought that fifteen minutes would be enough based on my previous implementations, but it was still insufficient for many students. As non-native speakers tend to take more time to understand the question, including requiring it to be read slowly and multiple times, it is important to plan sufficient time. Indeed, several students were pressed for time in a way that meant I could not ask all the probing follow-up questions that I would have liked, especially on the last question. Additionally, I suggest “priming” students for the type of question you will ask next during the exam to aid in their comprehension. For example, saying something like “this next question will be about Bayesian and frequentist probability interpretations” can help students follow the ensuing question more easily without needing it to be repeated.
@@
I also found the grading to be more difficult than I expected, particularly for partially correct answers. For example, students might have said an extreme result was more “surprising” but could not clearly articulate what this meant. Moreover, it was challenging to keep consistent grading expectations throughout the day when the first and last exam were close to eight hours apart. For example, I started to realize that one of the Central Limit Theorem follow-up questions was particularly confusing for students, so half-way though the day, I changed what I was looking for in this response. Since I could re-listen to the previous responses and adjust their grades accordingly, this was not a problem, but still it created some challenges for fairness. If you plan to give an oral exam, the more precise and granular the rubric you can make, the better.
@@
Perhaps the biggest limitation of this study was the small sample size ($n=9$), leading to low power. That being said, if we consider the data in Figure 2 and test $H_{0}:\rho\geq 0$ vs. $H_{1}:\rho<0$ where $\rho$ is the Pearson correlation coefficient, then the resulting $p$-value is 0.09, providing moderate evidence against the null hypothesis that stronger English skills lead to better oral exam performance. In any case, instead of these results being interpreted as conclusive support that oral exams do not favor students with strong English-speaking abilities, they should be taken as illustrative and an encouragement for other instructors to experiment with oral exams.
In summary, the test did not merely gauge a students’ English ability, but yielded a meaningful picture of their statistical understanding. Even while working in their second-language, I was impressed by many of the students ability to express complex statistical concepts. Moreover, the exam setting yielded the added benefit of being able to immediately encourage students who did particularly well, easing much of their stress as they left my office.
@@ Acknowledgments
I would like to thank Sunghwan Byun, Herle McGowan and Clay Williams for their helpful comments. I also thank the two anonymous referees and editor whose comments greatly improved the quality of the manuscript.
@@
Appendix
The following is a modified mastery rubric, adapted from Theobold, (2021). While the exam was primarily graded based on the explanation in Section 3.2, this rubric was used as a post hoc method to check the scores.
  • 4 points (excellent) – Outstanding ability to articulate the concepts covered in the exam, with comprehensive and thoughtful understanding of the content of the assessment. Few if if any errors occur. Any errors that do occur are minor and do not cloud the understanding articulated during the exam.
  • 3 points (good) – Ability to articulate the central concepts covered in the exam, but several errors may occur during the assessment. These errors, however, do not cloud the central concepts of the assessment.
  • 2 points (fair) – Limited ability to articulate the concepts covered in the exam, explanations contain inconsistencies and demonstrate limited understanding of the content
  • 1 point (poor) – Little to no articulation of concepts covered in the exam, numerous inconsistencies and errors occur and demonstrate little to no understanding of the content
  • 0 points (no credit) - Unable to demonstrate any understanding of concepts